40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 1 16:23:48 2018, Koji, Summary, Computers, Transition of the main NFS disk on chiara 
    Reply  Wed Aug 1 19:41:15 2018, gautam, Summary, Computers, RTCDS recovery, c1ioo changes CDS_overview.png
       Reply  Wed Aug 1 20:44:57 2018, gautam, Summary, Computers, c1omc model (re?)created CDS_OVERVIEW_withOMC.png
          Reply  Thu Aug 2 20:54:18 2018, gautam, Summary, Computers, c1omc model looks stable 
       Reply  Thu Aug 2 20:47:29 2018, gautam, Summary, Electronics, X Green "Mystery" solved 
          Reply  Thu Aug 2 23:09:25 2018, rana, Summary, Computers, X Green "Mystery" solved 
          Reply  Fri Aug 3 14:35:56 2018, gautam, Summary, Electronics, EX AUX electronics power restored 
    Reply  Mon Aug 6 09:42:10 2018, Koji, Summary, Computers, Transition of the main NFS disk on chiara 
Message ID: 14125     Entry time: Thu Aug 2 20:47:29 2018     In reply to: 14122     Reply to this: 14127   14128
Author: gautam 
Type: Summary 
Category: Electronics 
Subject: X Green "Mystery" solved 

I walked down to the X end and found that the entire AUX laser electronics rack isn't getting any power. There was no elog about this.

I couldn't find any free points in the power strip where I think all this stuff was plugged in so I'm going to hold off on resurrecting this until tomorrow when I'll work with Steve.

Quote:

The X arm green does not stay locked to the cavity - the alignment looks fine, and the green flashes are strong, but the lock does not hold. This shouldn't be directly connected to anything we did today since the Green PDH servo is entirely analog.

ELOG V3.1.3-